# FastTrackTr:Towards Fast Multi-Object Tracking with Transformers

### v2

@@ Abstract
[Uncaptioned image][Uncaptioned image]
Figure 1: A comparison of the FPS and HOTA of various versions of FastTrackTr and the most advanced transformer-based models on a single NVIDIA 4090. The numerical precision marked with diagonal lines in the right figure is FP32, while that without lines is FP16.
footnotetext: Liao Pan, Yang Feng, Wu Di, Yu Jinwen, Zhao Wenhui and Liu Bo are with Key Laboratory of Information Fusion, School of Automation, Northwestern Polytechnical University,Xi’an,China. (e-mile: liaopan@mail.nwpu.edu.cn,yangfeng@nwpu.edu.cn,wu_di821@mail. nwpu.edu.cn, yujinwen@mail.nwpu.edu.cn, zwh2024202513@mail.nwpu.edu.cn,503790475@qq.com
@@ 1 Introduction
Multi-object tracking (MOT) has long been a computer vision task of great interest to researchers and developers, with extensive application prospects across various industries. Currently, the mainstream tracking models are primarily online models, and the paradigm that has garnered widespread attention in recent years within online models remains tracking-by-detection (TBD). This paradigm, originating from SORT [1], has stood the test of time and demonstrates remarkable vitality. Its improved derivative methods [19, 7, 41] continue to be considered state-of-the-art (SOTA).
In recent years, some tracking methods based on transformer [34] have also attracted considerable attention. However, these methods [31, 20, 43, 39, 10, 42, 11] face a significant challenge: suboptimal inference speed. There are various reasons for this, such as the slow performance of the MOTR series, which relies on query-based tracking. The main reason is not the network structure, but issues related to hardware acceleration caused by the variable number of queries (detailed justification can be found in the appendix). Both our experiments and other works [22] have also demonstrated that the inference speed after lightening the network is still not very ideal. Regarding another recent method, MOTIP [11], while it addresses many of MOTR’s issues, its main problem lies in its complex network structure. Even after replacing MOTIP’s Deformable DETR [51] with RT-DETR [48], its inference speed remains unsatisfactory (results shown in Figure 1).
@@
Furthermore, it is noteworthy that models with a single decoder are capable of achieving end-to-end tracking without increasing the number of queries. This point is further discussed in the subsequent sections and the appendix of this document.
@@
    proposing FastTrackTr, a novel transformer-based MOT method that achieves fast inference speed while maintaining high accuracy.
@@ 2.2 Tracking-by-Detection
The TBD paradigm can be considered one of the simplest and most practical tracking methods. Since it essentially associates detected targets, this class of methods often yields satisfactory tracking accuracy as long as the detection results are not poor.
The initial SORT algorithm simply utilized the motion state of detected objects for association. Subsequent models like DeepSORT [36] and ByteTrack [46] improved tracking accuracy by enhancing detection precision and employing embedding-based Re-identification (Re-id) networks. Some approaches, like OC-SORT [3] and HybridSORT [41], improved upon SORT using purely kinematics-based methods. Later models such as MotionTrack [24], DeepOC-SORT [19], and SparseTrack [16] demonstrated that simultaneous improvements in both Re-id network utilization and kinematic models yield the most significant performance enhancements.
Currently, there are also algorithms that directly use networks for association. Besides the aforementioned PuTR [14], SUSHI [5] is a tracking method based on graph neural networks.
@@ 3.1 FastTrackTr Architecture
The overall structure of FastTrackTr is illustrated in Figure 2. In frame $T_{1}$, a standard decoder is used to initialize the initial historical information. Subsequently, in the following frames, a Historical Encoder is employed to initialize this historical-representative information. In the next frame, a Historical Decoder, which includes an additional cross-attention layer compared to normal decoder layers, processes this historical historical information. The tracking principle is very similar to that of JDT and FairMOT, utilizing an additional ID Embedding head to obtain the object’s appearance embedding, which is then used along with detection results for association and matching. It is worth noting that the backbone network is ResNet [13].
@@ 3.2 Learning Detection and ID Embedding
The second objective addresses a metric learning problem, specifically the learning of an embedding space where instances of the same identity are close to each other while those of different identities are far apart. Previous methods like JDE [35] and FairMOT [45] used one-hot labels. While this approach is simple and effective, it requires prior knowledge of the number of IDs in the training set to configure the output dimensions of the ID embedding head. This requirement obviously limits the model’s generalization capability and becomes impractical in dense scenarios, as it leads to excessively high embedding dimensions. Moreover, subsequent Transformer-based tracking models predominantly rely on queries for tracking, and the drawbacks of this method have been discussed earlier.
MOTIP [11] employs a predefined ID dictionary and classification methods for object association, whereas PuTR utilizes an association method based on embedding similarity. These two approaches are undoubtedly more generalizable. However, possibly because their decoders primarily process the trajectories of object movements when outputting embeddings, and since the queries used by FastTrackTr to generate embeddings contain more image features and positional information of the current target, these methods fail to enable FastTrackTr to achieve satisfactory convergence.
To address the aforementioned issues, we need to find a method more suitable for image features and with greater generalizability for generating ID embeddings. Therefore, we turn our attention to the training methods of ReID (Re-identification) models. By converting the generation of ID embeddings into a ReID problem, we only need to consider the distances and similarities between embeddings, thereby enhancing the model’s generalization capability.
For the above reasons and for better performance, in this paper, we adopt the more advanced circle loss [33] in this work. Specifically, the id embeddings of the same object between multiple frames is treated as positive pairs denoted as $s_{p}$ and between different objects as negative samples denoted as $s_{n}$, respectively. - $\alpha_{p}$ and $\alpha_{n}$ are the adaptive weighting factors for positive and negative samples, determined by the following equations:
@@ 3.3 Historical Decoder and Encoder
Specifically, each decoder module in DETR [4] consists of a sequential structure comprising a self-attention layer for the queries, a visual cross-attention layer, and a feed-forward neural network (FFN). Initially, the self-attention layer performs global modeling over the queries to enhance the model’s object detection performance. To further leverage historical information and improve tracking accuracy, we concatenate the final decoder output from frame $t-1$, denoted as $q^{t-1’}_{f}$, which is processed by the trajectory encoder, with the initial queries $q_{d}^{t}$ of frame $t$. This concatenation replaces the key and value in the self-attention layer while keeping the queries unchanged. The implementation can be formally expressed as follows:
@@
This mechanism allows each object query to adaptively capture features of the same object from images across consecutive frames, thereby achieving better object understanding. The queries with estimated historical information are then input into the inter-query self-attention layer for feature interaction between objects, and the visual cross-attention layer for collecting visual semantics. We use 6 decoder layers.
@@
The workflow of the historical encoder is illustrated in Figure 4. After generating the mask, we fuse the decoder output $q_{f}^{t}$ with the historical information $q_{f}^{t-1}$ through an attention mechanism, computed as follows:
$Q_{E}=\quad K_{E}=\operatorname{Linear}\left(q^{t-1}_{f}+q_{pos}\right),V_{E}=\operatorname{Linear}(q^{t}_{f})$ (8)
The attention calculation follows the same formula as previously used in the decoder, which is not repeated here. Finally, the fused features are processed by an additional network layer to obtain $q^{t^{\prime}}_{f}$. It is noted that for $t=1$, $q^{t-1}_{f}=q^{t}_{f}$.
The mask generation module in this method includes two modes: (1) In the training phase, the mask is determined by ground truth information, generated by checking if each object matches with the ground truth labels, and random sampling of the masks is performed to flip them, enhancing the model’s robustness; (2) In the inference phase, the mask is determined based on the confidence score of each object, with objects below a certain threshold being filtered out, while high-confidence objects are retained. This masking mechanism ensures that the model focuses on processing significant historical information, thereby improving the encoder’s capability to model the temporal dynamics of target trajectories.
@@ 4 Experiments
Table 1: Overall Latency comparison between FastTrackTr and existing methods on the Dancetrack [32] val. Methods highlighted in gray utilize the YOLOX [12] detector, with its processing time included. Similarly, the latency of our method accounts for its associated time consumption, demonstrating efficient performance with the chosen configuration. $\text{Dec}^{3}$ refers to a 3-layer decoder layer, while R18 refers to ResNet18.
Methods Image Res. Numerical Precision Overall Latency(ms) HOTA
DeepOC-SORT[19] $1440\times 800$ FP16 28.1ms (35.6fps) 58.5
SUSHI[5] $1440\times 800$ FP32/FP16 42.9ms (23.3fps) / 38.6ms (25.9fps) ///
PuTR[14] $1440\times 800$ FP32/FP16 30.5ms (32.8fps) / 28.2ms (35.5fps) 55.7
MOTR [43] $1333\times 800$ FP32/FP16 60.0ms (16.5fps) / 52.9ms (18.9fps) 51.3
MOTIP[11] $1333\times 800$ FP32/FP16 61.7ms (16.2fps) / 45.7ms (21.8fps) 59.1
FastTrackTr $1333\times 800$ FP32/FP16 40.2ms (24.8fps) / 30.9ms (32.4fps) 56.8
MOTIP(RT-DETR) $640\times 640$ FP32/FP16 32.1ms (31.2fps) / 24.5ms (40.8fps) 56.2
FastTrackTr $640\times 640$ FP32/FP16 23.6ms (42.3fps) / 19.6ms (51.0fps) 54.1
FastTrackTr-$\text{Dec}^{3}$ $640\times 640$ FP32/FP16 20.4ms (49.0fps) / 14.8ms (67.6fps) 51.2
FastTrackTr-R18-$\text{Dec}^{3}$ $640\times 640$ FP32/FP16 12.8ms (78.0fps) / 9.86ms (101.4fps) 47.4
@@ 4.1 Setting
Implementation Details: For inference speed, our FastTrackTr is built on RT-DETR [48], with ResNet50 [13] as the backbone unless specified otherwise. The model is implemented in PyTorch [23], and the training is conducted on 8 NVIDIA RTX 4090 GPUs. The sequence length for training is set to 8 unless otherwise noted.
All ablation experiments use an image size of $640\times 640$. For further testing of the best tracking accuracy of our method, we use an image size of $800\times 1333$ when comparing with SOTA methods. The Trajectory Encoder during training applies random mask transformations, with the probability related to the number of ground-truth objects. Given the number of ground-truth objects $n_{gt}$ and the number of queries $n_{q}$, the probability of a non-masked flip is $\min(0.5,2\cdot n_{gt}/n_{q})$, while for the rest, the probability is $\min(\max(0.1,n_{gt}/n_{q}),0.2)$. For all three datasets, $n_{q}$ is set to 300. Dancetrack and SportMOT are trained for 27 epochs, while MOT17 is trained for 80 epochs. The optimizer is AdwmW [17] with a learning rate of $10^{-4}$ and weight decay of $5.0\times 10^{-4}$. Due to GPU memory limitations, the batch size is set to 4 for smaller images and 1 for larger images. Lastly, for all datasets, the final output dimension of ID embedding is 256.
@@ 4.2 Main Result
Comprehensive Comparison of Speed and Accuracy: Table 1 highlights FastTrackTr’s superior latency performance compared to existing methods on the Dancetrack validation set. FastTrackTr consistently delivers faster processing speeds while maintaining competitive HOTA scores, demonstrating efficient real-time capabilities. Additionally, its performance scales effectively with reduced image resolutions, further enhancing its applicability. Advanced variants of FastTrackTr achieve even higher frame rates with only minimal compromises in tracking accuracy. These advantages make FastTrackTr highly effective for practical multi-object tracking applications.
@@
Comparison of State-Of-The-Art: To better demonstrate the superiority of our FastTrackTr, we employed nearly all available data augmentations and used larger image sizes. Tables 2, 3, and 4 respectively showcase the maximum tracking performance of FastTrackTr on DanceTrack, SportsMOT, and MOT17 benchmark. Similar to many previous works [49, 47, 10, 43], we utilized the CrowdHuman [28] dataset for data augmentation on MOT17 and DanceTrack. It is important to note that although our method still does not outperform CNN-based models on the MOT17 benchmark, the underlying reason remains the well-discussed issue: for transformer models, the amount of data in the MOT17 dataset is too limited. Nevertheless, our method has surpassed all previous transformer-based approaches. This indicates that our method performs better even when the dataset images do not allow for sufficient training of the model, suggesting that our approach has potential for practical application in real-world settings, rather than being confined mostly to academic research like previous transformer methods, with few exceptions.
Methods HOTA* DetA AssA MOTA IDF1
FairMOT [45] 49.3 70.2 34.7 86.4 53.5
QDTrack [9] 60.4 77.5 47.2 90.1 62.3
ByteTrack [46] 62.1 76.5 50.5 93.4 69.1
OC-SORT [3] 68.1 84.8 54.8 93.4 68.0
MeMOTR [10] 68.8 82.0 57.8 90.2 69.9
FastTrackTr(ours) 70.1 83.9 58.6 94.0 72.0
@@
FastTrackTr(our) 62.4 62.8 63.0 76.7 77.2

### v3

# FastTrackTr:Towards Real-time Multi-Object Tracking with Transformers

@@ Abstract
Figure 1: A comparison of the FPS, HOTA and MOTA of various versions of FastTrackTr and the most advanced transformer-based models on a single NVIDIA 4090. For the left image, the horizontal axis is FPS, the vertical axis is HOTA, and the radius of the shape is MOTA.
@@ 1 Introduction
Multi-object tracking (MOT) has long been a computer vision task of great interest to researchers and developers, with extensive application prospects across various industries. Currently, the mainstream tracking models are primarily online models, and the paradigm that has garnered widespread attention in recent years within online models remains tracking-by-detection (TBD). This paradigm, originating from SORT [2], has stood the test of time and demonstrates remarkable vitality. Its improved derivative methods [21, 8, 43, 1] continue to be considered state-of-the-art (SOTA).
In recent years, some tracking methods based on transformer [36] have also attracted considerable attention. However, these methods [33, 22, 45, 41, 11, 44, 12] face a significant challenge: suboptimal inference speed. There are various reasons for this, such as the slow performance of the MOTR series, which relies on query-based tracking. The main reason is not the network structure, but issues related to hardware acceleration caused by the variable number of queries (detailed justification can be found in the appendix). Both our experiments and other works [24] have also demonstrated that the inference speed after lightening the network is still not very ideal. Regarding another recent method, MOTIP [12], while it addresses many of MOTR’s issues, its main problem lies in its complex network structure. Even after replacing MOTIP’s Deformable DETR [53] with RT-DETR [50], its inference speed remains unsatisfactory (results shown in Figure 1). And, during its implementation, some of the data is dynamic, which makes it relatively difficult to accelerate using TensorRT.
@@
Furthermore, the deterministic data architecture inherent in FastTrackTr enables seamless TensorRT acceleration, achieving genuine real-time performance. This computational efficiency remains largely unattainable for contemporary transformer-based MOT methods due to their inherent dynamic computation patterns.
@@
    proposing FastTrackTr, a novel transformer-based MOT method that achieves real-time inference speed while maintaining high accuracy.
@@ 2.2 Tracking-by-Detection
The TBD paradigm can be considered one of the simplest and most practical tracking methods. Since it essentially associates detected targets, this class of methods often yields satisfactory tracking accuracy as long as the detection results are not poor. The initial SORT algorithm simply utilized the motion state of detected objects for association. Subsequent models like DeepSORT [38] and ByteTrack [48] improved tracking accuracy by enhancing detection precision and employing embedding-based Re-identification (Re-id) networks. Some approaches, like OC-SORT [4] and HybridSORT [43], improved upon SORT using purely kinematics-based methods. Later models such as MotionTrack [26], DeepOC-SORT [21], and SparseTrack [18] demonstrated that simultaneous improvements in both Re-id network utilization and kinematic models yield the most significant performance enhancements. Currently, there are also algorithms that directly use networks for association. Besides the aforementioned PuTR [16], SUSHI [6] is a tracking method based on graph neural networks.
@@ 3.1 FastTrackTr Architecture
The overall structure of FastTrackTr is illustrated in Figure 2. In frame $T_{1}$, a standard decoder is used to initialize the initial historical information. Subsequently, in the following frames, a Historical Encoder is employed to initialize this historical-representative information. In the next frame, a Historical Decoder, which includes a historical cross-attention layer compared to normal decoder layers, processes this historical information. The tracking principle is very similar to that of JDT and FairMOT, utilizing an additional ID Embedding head to obtain the object’s appearance embedding, which is then used along with detection results for association and matching. It is worth noting that the backbone network is ResNet [14].
@@ 3.2 Learning Detection and ID Embedding
The second objective addresses a metric learning problem, specifically the learning of an embedding space where instances of the same identity are close to each other while those of different identities are far apart. Previous JDT methods (e.g., JDE [37], FairMOT [47]) rely on one-hot labels to construct ID embeddings, necessitating predefined ID quantities to configure embedding dimensions. This design introduces two critical limitations: (1) explosive dimensionality growth in dense scenes and (2) severely restricted generalization due to dataset-specific ID dependencies. Subsequent Transformer-based trackers inherit these constraints through their query mechanisms. While recent approaches like MOTIP [12] (using ID dictionaries) and PuTR (leveraging embedding similarity) improve generalizability, their trajectory decoders overemphasize motion patterns, conflicting with FastTrackTr’s queries that encode richer spatial features and positional priors. This mismatch hinders effective convergence. To resolve these issues, we reformulate ID embedding learning as a re-identification (ReID) problem, decoupling embedding dimensions from ID quantities and prioritizing feature distance/similarity optimization. This paradigm shift enhances cross-scene adaptability and better aligns with image-driven tracking frameworks.
For the above reasons and for better performance, in this paper, we adopt the more advanced circle loss [35] in this work. Specifically, by utilizing the dot product similarity, the id embeddings of the same object across multiple frames are regarded as positive pairs indicated as $s_{p}$, while those between distinct objects are treated as negative samples represented by $s_{n}$. - $\alpha_{p}$ and $\alpha_{n}$ are the adaptive weighting factors for positive and negative samples, determined by the following equations:
@@ 3.3 Historical Decoder and Encoder
Specifically, each decoder module in DETR [5] consists of a sequential structure comprising a self-attention layer for the queries, a visual cross-attention layer, and a feed-forward neural network (FFN). Initially, the self-attention layer performs global modeling over the queries to enhance the model’s object detection performance. To further leverage historical information and improve tracking accuracy, we concatenate the Historical Encoder output from frame $t-1$, denoted as $q^{t-1’}_{f}$, which is processed by the trajectory encoder, with the initial queries $q_{d}^{t}$ of frame $t$. This concatenation replaces the key and value in the self-attention layer while keeping the queries unchanged. The implementation can be formally expressed as follows:
@@
This mechanism allows each object query to adaptively capture features of the same object from images across consecutive frames, thereby achieving better object understanding. The queries with estimated historical information are then input into the inter-query self-attention layer for feature interaction between objects, and the visual cross-attention layer for collecting visual semantics.
@@
The workflow of the historical encoder is illustrated in Figure 4. After generating the mask, we fuse the decoder output $q_{f}^{t}$ with the historical information $q_{f}^{t-1^{\prime}}$ through an attention mechanism, computed as follows:
$Q_{E}=K_{E}=\operatorname{Linear}(q^{t}_{f}+q_{pos}),V_{E}=\operatorname{Linear}(q^{t-1^{\prime}}_{f})$ (8)
where $q_{pos}$ represents a set of learnable positional encodings and the attention weights are computed via scaled dot-product as previously used in the decoder, followed by masking to suppress less relevant historical elements. Finally, the fused features are then processed by a feed-forward network to obtain $q^{t^{\prime}}_{f}$. For $t=1$, we initialize $q^{t-1^{\prime}}_{f}$ as $q^{t}_{f}$ due to the absence of prior history.
To prioritize significant historical information, we design a dynamic masking strategy with two operational modes: (1) In the training phase, the mask is determined by ground truth information, generated by checking if each object matches with the ground truth labels, and random sampling of the masks is performed to flip them, enhancing the model’s robustness; (2) In the inference phase, the mask is determined based on the confidence score of each object, with objects below a certain threshold being filtered out, while high-confidence objects are retained. The mask adopts [36]’s softmax suppression via large negative biases to irrelevant entries. It strengthens the encoder’s capability to model temporal dependencies while suppressing error propagation from noisy historical inputs.
@@ 4 Experiments
Table 1: Overall Latency comparison between FastTrackTr and existing methods on the Dancetrack [34] val. * denotes that these methods employ the same YOLOX [13] detector, and the latency of this detector is accounted for in this table. Similarly, the latency of our method accounts for its associated time consumption, and these are highlighted in color. $\text{Dec}^{3}$ refers to a 3-layer decoder, while R18 refers to ResNet18. Unless otherwise specified, all Transformer-based models have ResNet50 for Backbone and 6 layers for decoder.
Methods Image Res. Numerical Precision Overall Latency(ms) HOTA Overall Latency with TensorRT FP16(ms)
Deep OC-SORT*[21] 1333×800 FP16 36.6ms (27.8fps) 58.5 13.5ms(74.1fps)
SUSHI*[6] 1333×800 FP32/FP16 42.7ms (23.4fps) / 38.0ms (26.3fps) 58.7 -
PuTR*[16] 1333×800 FP32/FP16 29.9ms (33.4fps) / 27.7ms (36.1fps) 54.5 -
MOTR[45] 1333×800 FP32/FP16 60.0ms (16.5fps) / 52.9ms (18.9fps) 51.3 -
MOTIP[12] 1333×800 FP32/FP16 61.7ms (16.2fps) / 45.7ms (21.8fps) 59.1 -
FastTrackTr 1333×800 FP32/FP16 40.2ms (24.8fps) / 30.9ms (32.4fps) 56.8 11.6ms(86.6fps)
MOTIP(RT-DETR) 640×640 FP32/FP16 32.1ms (31.2fps) / 24.5ms (40.8fps) 56.2 -
FastTrackTr 640×640 FP32/FP16 23.6ms (42.3fps) / 19.6ms (51.0fps) 54.1 6.0ms(166.4fps)
FastTrackTr-Dec3 640×640 FP32/FP16 20.4ms (49.0fps) / 14.8ms (67.6fps) 51.2 4.9ms(203.7fps)
FastTrackTr-R18-Dec3 640×640 FP32/FP16 12.8ms (78.0fps) / 9.86ms (101.4fps) 47.4 4.1ms(243.4fps)
@@ 4.1 Setting
Implementation Details: For inference speed, our FastTrackTr is built on RT-DETR [50], with ResNet50 [14] as the backbone unless specified otherwise. The model is implemented in PyTorch [25], and the training is conducted on 8 NVIDIA RTX 4090 GPUs. The sequence length for training is set to 8 unless otherwise noted. All ablation experiments use an image size of $640\times 640$. For further testing of the best tracking accuracy of our method, we use an image size of $800\times 1333$ when comparing with SOTA methods. The Historical Encoder during training applies random mask transformations, with the probability related to the number of ground-truth objects. Given the number of ground-truth objects $n_{gt}$ and the number of queries $n_{q}$, the probability of a non-masked flip is $\min(0.5,2\cdot n_{gt}/n_{q})$, while for the rest, the probability is $\min(\max(0.1,n_{gt}/n_{q}),0.2)$. For all three datasets, $n_{q}$ is set to 300. Dancetrack and SportMOT are trained for 27 epochs, while MOT17 is trained for 80 epochs. The optimizer is AdwmW [19] with a learning rate of $10^{-4}$ and weight decay of $5.0\times 10^{-4}$. Due to GPU memory limitations, the batch size is set to 4 for smaller images and 1 for larger images. Lastly, for all datasets, the final output dimension of ID embedding is 256. All models are implemented in PyTorch by default unless explicitly stated otherwise. The complete codebase remains exclusively Python throughout our experiments.
@@ 4.2 Main Result
Comprehensive Comparison of Speed and Accuracy: In RT-DETR, a standardized benchmark has been established for real-time detectors. Within this benchmark, all models undergo acceleration via TensorRT, with computational latency measurements exclusively encompassing inference and post-processing durations. This paper adopts the same configuration while substituting the detection dataset COCO2017 [15] with the MOT dataset Dancetrack . However, this substitution reveals a critical implementation challenge: most transformer-based MOT methods [33, 22, 45, 41, 11, 44, 12, 16] developed in PyTorch incorporate dynamic computational graphs during inference, presumably to reduce computational overhead or simplify implementation. These dynamic elements create significant compatibility issues (specifically non-fixed tensor shapes and conditional branching) when converting models to TensorRT - challenges that remain unresolved in our current implementation. To ensure fair comparison, Table 1 comprehensively presents both speed and accuracy metrics across tracking methods using various numerical precisions in PyTorch. Notably, for our proposed models and the TBD method, only the network benefits from TensorRT acceleration, while the association module retains its original Python implementation.
@@
Comparison of State-Of-The-Art: To better demonstrate the superiority of our FastTrackTr, we employed nearly all available data augmentations and used larger image sizes. Tables 2, 4, and 3 respectively showcase the maximum tracking performance of FastTrackTr on DanceTrack, SportsMOT, and MOT17 benchmark. Similar to many previous works [51, 49, 11, 45], we utilized the CrowdHuman [30] dataset for data augmentation on MOT17 and DanceTrack. It is important to note that although our method still does not outperform CNN-based models on the MOT17 benchmark, the underlying reason remains the well-discussed issue: for transformer models, the amount of data in the MOT17 dataset is too limited. Nevertheless, our method has surpassed all previous transformer-based approaches. This indicates that our method performs better even when the dataset images do not allow for sufficient training of the model, suggesting that our approach has potential for practical application in real-world settings, rather than being confined mostly to academic research like previous transformer methods, with few exceptions.
@@
FastTrackTr(ours) 62.4 62.8 63.0 76.7 77.2
Methods HOTA* DetA AssA MOTA IDF1
FairMOT [47] 49.3 70.2 34.7 86.4 53.5
QDTrack [10] 60.4 77.5 47.2 90.1 62.3
ByteTrack [48] 62.1 76.5 50.5 93.4 69.1
OC-SORT [4] 68.1 84.8 54.8 93.4 68.0
MeMOTR [11] 68.8 82.0 57.8 90.2 69.9
FastTrackTr(ours) 70.1 83.9 58.6 94.0 72.0